latanya make a 17 minute phone call from her hotel in france and was chargered 17 euro. at the time 1 $ was 0.83 euro. find the cost per minute of the call in dollars
Answer by rfer(16322)

1/0.83=x/17
0.83x=17
x=17/0.83
x=21.69